(a) Calculate the percent ionization of 0.00640 M hypobromous acid (Ka = 2.5e-09) % ionization = (b) Calculate the percent io(b) Calculate the percent ionization of 0.00640 M hypobromous acid in a solution containing 0.0790 M sodium hypobromite.
Answer #1

Blution Given or data .. 1 hypobromous acid Ka 20.00 6uo M a 2.5 410-09 Brow -> Bro +H+ Initial 0.006uo . x x Fruilibrium 0-c[ro] (the ot] Sasklon 6.00CRO-X U = 0.00640 x Svaly CHOO 640 - 2.5810- = gel - 0.0064.0 . 1.6460-1l M. * = 1.6xre M = 2t] =(13) typobromous acid obromite = 0.00640 M = 0.0790 M 4-toBr + H20 - Initial oroobuo change - Bro 0.079 +H30- o 0.079 +2 -0.0% ionization =1t307 HoBryxion a 3-15 X10 x100 3. 15 X10